1.The half-life of chromium-51 is 28 days. If the initial amount of chromium-51 is 133 mCi, how much chromium-51 is left in the body after 84 days?


1
Expert's answer
2021-05-26T07:07:23-0400

N(t)=Nо(1/2)^t/1/2t

N(t)-remaining amount of substance

N(о)-Initial amount of the substance

N(t)=1/2*133mCi^84/1/2*28

N(t)=133*1/2^6

N(t)=133*1/64

N(t)=133/64

N(t)=2,078125 ~~ 2,08